OpenStudy (anonymous):

If there is an arithmatic sequence where a7= 12 and a19=-36 how do you find the general nh term formula for the 32nd term?

OpenStudy (anonymous):

a7 = a1 + 6*d a19 = a1 + 18*d a32 = a1 + 31d i) 12 = a1 + 6d ii) -36 = a1 + 18d iii) a32 = a1 + 31d i - ii: 12 - (-36) = a1 + 6d - (a1 + 18d) 48 = - 12d iv) d = -4 iv -> i: 12 = a1 + 6*(-4) 12 = a1 - 24 a1 = 36 iii) a32 = (36) + 31*(-4) a32 = -88

OpenStudy (anonymous):

thank you vitor023